

********************************************************************
UARS WEEKLY STATUS REPORT - 22 MAR 2002
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, taking solar data on all
available orbital tracking periods.
***********
SUSIM
***********
SUSIM performed its daily mid-resolution and MgII solar scans during
periods of predicted UARS telemetry.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
A detailed review of the PEM CECIL PROCs has been completed. All 51 PEM
CECIL PROCs have been approved. PEM thanks Greg Haun for his assistance with
the CECIL conversion.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered.
PEM awaits support of the UARS Traceability Mission for the TIMED mission.
Initial meetings with a GUVI team representative have highlighted the need
for knowledge of conjunction periods. Investigation is underway into
identifying times when the PEM spectrometers and the GUVI view plane are
sampling from the same volume of the atmosphere. Orbital investigations
indicate that the region of time around April 28 looks promising. Initial
analysis of conjunction times indicate that the terminator is close to the
conjunction point. This effects the quality of the GUVI retrieval, which
works best in either total sunlight or total darkness.
Magnetic conditions were investigated for this time frame. Data from
April 19 shows the best magnetic conditions; however the terminator location
is in the worst location for GUVI. A second conjunction on April 23 does
not have great magnetic conditions and the TIMED orbit makes the GUVI retrieval
suspect. The magnetic conditions on April are not favorable; however, GUVI
is able to invert to obtain a solution. Additional investigation is currently
being conducted.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
WINDII was activated on March 13, in its "warm" state; that is,
with active thermal control. This will allow the zero wind
phase to be traced back to the beginning of the mission, for
the validation with TIDI on TIMED. Operations are planned
to continue to April 8. Coincidentally this run covers the
period of the "springtime" transition which is of interest
to both WINDII and TIDI. With two spacecraft the associated
planetary waves will be more accurately characterized than is
possible with just one instrument.
***********
MPG
***********
The MPG attended the ECLIPSE Operational Readiness
Review on 21 March 2002.
The MPG continues to serve as the point-of-contact
for UARS instrument operations to support re-engineering
of the UARS Mission Operations Concept to develop
capabilities for the Science Traceability Mission.
Daily processing of instrument Daily Activity Plans
continues with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 09 March 2002 (Orbit 57364, GMT Day 068) through 15 March
2002 (Orbit 57468, GMT Day 074).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
SUSIM 070/1338 57402 Power ON (Part 1)
SUSIM 071/1020 57415 Power ON complete
WINDII 072/0004 57424 Power ON
SOLSTICE 072/0547 57427 Day Only
HRDI 072/2035 57436 Power ON - Low Power
HRDI 072/2118 57437 Power ON - Low Power
HRDI 073/0139 57440 Power ON Day Only
PEM 073/1239 57446 Power ON day only
HALOE 074/0038 57454 Power ON part 1
HALOE 074/0914 57459 Power ON complete
The beta angle ranged between the angles of 2.72 degrees to 23.92
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 54 degrees on 27 March 2002 (DOY 086).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 25.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 1.25 degrees C. Additional battery
performance data for each battery on 09 March 2002 and 15 March 2002
are provided in Appendix D.
The clock error ranged from -11.6 to 7.4 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 2
HRDI = 3
HALOE = 2
SOLSTICE = 1
PEM = 1
SUSIM = 2
WINDII = 1
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 8
Three (3) Anomaly Report remained OPEN at the end of this report
period (see Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 01/333 - 2287.496855
Transponder B: 02/043 - 2287.498879
GROUND SYSTEM OPERATION
Eight (8) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
Eclipse testing remained the focal point this past week on UARS. The
integration of the new system has reached its final month and only a
few weeks are left before the April 1st deadline. The FOT continues
to test all of the Eclipse functionality against the UTTS and against
the spacecraft. Some errors have been found but the Raytheon software
development team has done an excellent job of turning around the fixes
for these problems.
MEETINGS
The FOT attended a UARS Eclipse Working Meeting on Thursday, March 14,
2002.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 13 April 2002 (DOY 103). This
maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 15 March 2002
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
57375 068/1856 02-077 OBC Clk Adj fr -20.6 to 0.0 msec/day
57402 070/1338 02-078 SUSIM Power ON (Part 1)
57415 071/1020 02-078 SUSIM Power ON complete
57424 072/0004 02-079 WINDII Power ON
57427 072/0547 02-080 SOLSTICE Day only ops due to low power
57436 072/2035 02-081 HRDI Power ON - Low Power
57437 072/2118 02-081 HRDI Power ON - Low Power
57440 073/0139 02-082 HRDI Power ON day only
57446 073/1239 02-083 PEM Power ON day only
57454 074/0038 02-084 HALOE Power ON part 1
57459 074/0914 02-085 HALOE Power ON complete
57467 074/2215 02-086 OBC Ground Overlay Fix
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 15 March 2002
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
____________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 15 March 2002
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
57375 068/1856 Clk Adj fr -20.6 to 0.0 msec/day
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 15 March 2002
09 March 2002 - GMT Day 068 Beta = 02.72 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -44.8/-56.0 -3.77/-4.99 +0.00/-0.00 10.72 0.000 0.00
2 00.0/-123.2 +4.81/+4.15 +27.2/-12.8 25.76 1.020 20.27
3 +16.8/-50.4 +3.50/+2.85 +26.8/-12.8 25.60 1.020 19.92
15 March 2002 - GMT Day 074 Beta = 23.92 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- ---------- ---------- ----- ------- ------
1 -44.8/-50.4 -3.77/-4.69 +0.0/ -0.0 10.72 0.000 0.00
2 00.0/-72.8 +5.14/+4.48 +28.0/-14.0 26.08 1.020 20.16
3 +22.4/-22.4 +3.82/+3.17 +26.8/-13.6 25.92 1.020 19.81
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---- ---------
-no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 15 March 2002
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-006 FOT 57364 068/0046 N/A PMON #1 Tripped (Low Pwr)
02-007 FOT 57450 073/1747 N/A AP-7 Crashed
02-008 FOT 57452 073/2117 N/A AP-1 Crashed
02-009 FOT 57450 073/2117 N/A AP-7 Crashed
02-010 FOT N/A N/A Prime Pack (PE-54) Crash
02-011 FOT 57466 074/2033 N/A AP-7 Disc Drive Failure
02-012 FOT 57466 074/2033 N/A AP-1 Hang-up
02-013 FOT 57467 074/2116 N/A AP-1 Hang-up
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-006 FOT 57364 068/0046 N/A PMON #1 Tripped (Low Pwr)
02-007 FOT 57450 073/1747 N/A AP-7 Crashed
02-008 FOT 57452 073/2117 N/A AP-1 Crashed
02-009 FOT 57450 073/2117 N/A AP-7 Crashed
02-010 FOT N/A N/A Prime Pack (PE-54) Crash
02-012 FOT 57466 074/2033 N/A AP-1 Hang-up
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-004 SOLSTICE 56883 036/0050 N/A Solstice in Wrong Mode
02-011 FOT 57466 074/2033 N/A AP-7 Disc Drive Failure
02-013 FOT 57467 074/2116 N/A AP-1 Hang-up